5e+4f=9
4e+5f=9
how do u solve this in linear equation form,in substitution, and graphing

Answer by MORUPHOSUOLALE(24)



SOLUTIION
___________

5e+4f=9 -eq--1
4e+5f=9 -eq--2
This is a simulateneous equation.
By subtituion method
from -eq 1
f= (9-5e)/4

plug in (9-5e)/4 for f in eq-2
4e+5(9-5e)/4 = 9

4e + [{45-25e}/4]=9
16e+45-25e=36 BY multiplying through by 4 and open the bracket.
-9e=36-45
-9e=-9
e=9/9
=1
substitue e=1 for e in any of the euation to get the value of f
4e+5f=9-eg---2 above
4(1}+5f=9
4+5f=9
5f=9-4
5f=5
f=5/5
f=1
SECONDLY THE GRAPH OF
e-1=0 means e=1
so graph of F(e)= e-1 is a straight line graph
similarly the graph of f-1=0 means f=1 is also a straight line graph
My computer is not comapatible with drawing graph presently if not that i will be more than glad to show you how to draw the graph.
But according to your problem, we can generalize that e=f=1
Which means the value of e and f is the same.
They are both a straight line graphs.
MORUPH .A.OSUOLALE
CENTRAL PIEDMONT COLLEGE
CHARLOTTE,NC